# A Global Minimum Tax for Large Firms Only:

### v5

@@ 2.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its exogenous pre-tax profits $\pi$ from their non-haven affiliate to each of the tax haven affiliates $h$. Firms incur transaction costs when shifting profits. These costs can either be thought of as concealment costs, which have to be incurred to hide the profit shifting from the non-haven country’s tax authorities, or as the expected fine that is to be paid when profit-shifting is detected and sanctioned. We assume that the costs of shifting $\theta_{h}\pi$ of profits to each haven $h$, given by $C_{h}(\cdot)$, are convex and, for analytical simplicity, quadratic in the level of profit-shifting, $\theta_{h}\pi$. Higher total (true) profits make it easier, however, to hide a given amount of profit-shifting. [^13] This specification is widely used in the literature on profit shifting (e.g., Hines and Rice,, 1994; Huizinga and Laeven,, 2008; Suárez Serrato,, 2018). Therefore, the shifting costs for a firm with profits $\pi$ to a haven $h$ are $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. The exogenous parameter $\delta$ captures the ease with which profits can be shifted across countries. It incorporates any efforts that countries take to prevent profit-shifting. [^14] In Hindriks and Nishimura, (2022), this cost parameter is endogenously chosen by governments, either cooperatively or non-cooperatively.
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the tax haven countries, a MNE maximizes its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\Pi}$, by choosing the share of profit shifting to each haven:
@@
The MNEs’ optimal level profit shifting to each tax haven, $\theta_{h}$ is then:
@@
which is independent of $\pi$ and therefore holds for all MNEs simultaneously. [^15] In principle, our analysis can incorporate $\theta<0$, which occurs if $t_{n}<t_{h}$. However, this case will never arise in the tax-competition equilibrium we will see shortly.
The optimal $\theta$ defined by (2) allows us to incorporate the MNEs’ adjustment to tax differentials in a very simple and compact form. It implies that MNEs of different size will respond to international tax differentials with the same tax base elasticity. Substituting (2) into (1) and using the symmetry of tax havens, the tax base of a MNE with profits $\pi$ in the non-haven country is $TB_{n}=(1-\sum_{h=1}^{H}\theta_{h})\pi=[1-H(t_{n}-t_{h})/\delta]\pi$. This yields a tax base elasticity of an individual MNE in the non-haven country such that $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, which is independent of $\pi$.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^16] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al., (2022) also find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different size. Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark. As a consequence, there is no reason per-se in our model that countries differentiate tax rates between MNEs of different size. However, as we will see below, a split tax regime can nevertheless occur in equilibrium as a result of introducing a GMT for large firms only.
@@ 2.3 Governments’ tax setting choices
Before examining the effects of GMT, we first solve the unconstrained optimization problem in the absence of a GMT. The non-haven country $n$ maximizes a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, where all MNE owners reside exclusively in the non-haven country. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal valuation of public funds.
The welfare premium on tax revenues, relative to private income, is a standard assumption in international taxation, which motivates positive tax rates in equilibrium. It can be given two alternative interpretations. The first is a redistributive (or political-economy) motive, when the government uses the corporate tax to redistribute income from capital owners to other households in the economy, which possess no earned income. The other interpretation is that corporate tax revenue is used to reduce other distortive taxes, and the (exogenous) deadweight loss associated with them. [^17] The latter argument assumes that there is a fixed excess burden of taxation in a country’s tax system (see e.g., Keen and Lahiri,, 1998). This is a plausible assumption in our setting, as we focus on the corporation tax, which accounts for less than 10% of total tax revenue in all developed countries. Either setting leads to a welfare gain for the non-haven country from levying a positive corporation tax on its MNEs. Finally, for $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues. This special case underlies, for example, the analyses in Hindriks and Nishimura, (2022) or Janeba and Schjelderup, (2023).
The government of country $n$ therefore maximizes
[…]
@@
$\qquad\qquad+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
[…]
Each of the tax haven countries $h$ only cares about its tax revenues (multiplied by the same marginal value of public funds, $\lambda$, as for country $n$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
[…]
@@
Solving the first-order conditions of the two countries yields best responses: [^18] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
@@
Solving for the optimal tax rates in the unconstrained benchmark (“Regime 0”) then gives
@@
[…]
A few comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the marginal value of public funds, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the valuation of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. Finally, each of the havens without an independent tax base will choose one-half of the tax rate of the non-haven. [^19] Note the important difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. This contrast arises from different specifications of profit-shifting costs. Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. This winner-takes-all situation among the havens drives their equilibrium tax rates to zero. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs. Hence each MNEs has an incentive to diversify its allocation of shifted profits across all havens. As we discuss in our calibration in Section 4, a setting with positive tax rates in haven countries in the unconstrained tax equilibrium is more in line with the empirical evidence.
We can also derive the share of shifted profits in this unconstrained tax equilibrium. This is given by:
@@
Hence, when the non-haven country maximizes tax revenues ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium. For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
@@ 2.4 Tax regimes
With a GMT in place, there is a first stage of the game in which each country chooses to commit or not to a single tax rate, which can either be the GMT rate or a tax rate above the GMT. If a country does not commit to a single tax rate, then it will choose to split tax rates in the second stage, and set different tax rates for large and small MNEs. MNEs with profits equal to or greater than the threshold $\pi_{M}$, i.e., $\pi\in[\pi_{M},\infty)$, are subject to the GMT requirement (large MNEs), whereas those with $\pi\in[\underline{\pi},\pi_{M})$ are not (small MNEs). We denote by $\phi$ the share of total profits earned by large MNEs, or the coverage rate of the GMT, which is exogenously given:
[…]
@@
Table 1: Possible regimes of the non-cooperative tax game
Table 1 shows that this sequence yields three potential choices for each of the two sets of countries. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these potential regimes, if the non-haven sets a single GMT rate, the havens would choose a single tax that is at the same level or even higher. Hence tax revenues in each haven country would be zero. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven country for small MNEs. Alternatively, if the haven sets a single non-GMT or a single GMT tax rate, the non-haven will never want to split its tax rate, because the tax base elasticity of its MNEs is independent of profits. Hence only five possible tax regimes remain.
In Regime 0, the GMT rate is not binding for any country; this is the case of unconstrained tax competition analyzed in the previous section. As the GMT rate is continuously increased, the tax competition equilibrium passes through four other regimes. In the following, we first state the full characterization of the sub-game perfect Nash equilibrium of the three-stage tax competition game, before commenting on each of these regimes.
@@
Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal valuation of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterized by the following four regimes, with all symmetric haven countries choosing the same tax schedule.
@@
    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\pi_{M},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\pi_{M},\infty).$
@@
Proof: See Appendix A.1.
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates. Our discussion below will focus on tax revenue effects. This is strictly correct only for the tax havens. It is also possible for the non-haven, however, because the marginal value of public funds, $\lambda$, is constant in our analysis, and it always exceeds the valuation of private income. All formal proofs are relegated to the Appendix.
@@
In Regime 1, where $t_{M}$ rises above the haven’s unconstrained tax rate $t_{h}^{0}$, the haven is forced to set $t_{M}$ for large MNEs. By contrast, the non-haven’s unconstrained rate is still above the GMT. As the non-haven is not bound by the GMT, it will continue to choose a single tax rate above the GMT for all MNEs. For the tax havens, the choice is therefore whether to apply the GMT also to small MNEs, or use a split tax schedule. By committing to a single GMT rate for all MNEs, the havens can induce the non-haven to set a higher uniform tax rate than in the unconstrained equilibrium. This argument is strongest when tax havens collectively decide on whether to commit to a single tax rate, as we assume in our main model. [^20] This is seen from the non-haven’s best response function (5), which is increasing in $t_{h}$. However, as we show in Appendix A.2, the same qualitative argument also holds for decentralised decision-making of tax havens. Havens therefore trade off the gains from inducing a higher uniform tax rate in the non-haven against the revenue losses from not undercutting the non-haven’s tax rate even more for small MNEs. In Regime 1, the first effect dominates and the havens find it profitable to commit to the GMT for all MNEs. As a result, all countries choose higher uniform tax rates in Regime 1, as compared to the unconstrained equilibrium without a GMT.
In Regime 2, where $t_{M}$ rises above the regime-switching rate $t_{M}^{1}$, the non-haven is still not bound by the GMT, and will continue to set a single tax rate above the GMT for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. As the GMT is now higher, there is less to gain from setting a uniform GMT rate for all MNEs, because the tax differential and thus shifted profits are reduced. [^21] This is seen from Regime 2 of Proposition 1, where the tax differential, $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases in $t_{M}$. At the same time, given the higher GMT rate, there is more to gain for the havens from undercutting the non-haven’s tax rate for small MNEs. In Regime 2, therefore, the havens will find it optimal to split their tax rate and set $t_{M}$ only for the large MNEs, but choose a lower tax rate for small MNEs. [^22] In practice, this split is done by choosing the lower tax rate for small MNEs as the general corporate tax rate, but topping up this tax rate for large MNEs. See footnote 5 in the introduction. Johannesen, (2022) has shown formally that it is always rational for tax havens to use the QDMTT up to the total tax level of the GMT, rather than leaving the top-up taxation to the home countries of the MNEs. Note, finally, that the upper boundary of Regime 2, $t_{M}^{2}$, is rising in the GMT coverage rate, $\phi$. This is because the non-haven country faces less competition from the havens (i.e., higher tax rates $t_{h}$) for large firms. If large MNEs make up a large part of the corporate tax base, the non-haven country will therefore be able to set a uniform tax rate above $t_{M}$ for a wider range of GMT rates.
@@
In Regime 3, where $t_{M}$ rises above the regime-switching rate $t_{M}^{2}$, the non-haven will no longer charge a tax rate above the GMT level, but it is bound by the GMT tax rate $t_{M}$ for large firms. By committing to set $t_{M}$ for all MNEs, it induces the haven countries to charge a higher tax rate for small MNEs. This argument is analogous to the one that applied to the tax havens in Regime 1. In equilibrium, therefore, the non-haven sets $t_{M}$ for all firms, rather than splitting its tax rate, and the havens respond by charging $t_{M}/2$ for small firms.
@@
Finally, in Regime 4, where $t_{M}>t_{M}^{3}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set the GMT rate for all firms. [^23] This is seen from Regime 3 of Proposition 1, where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases in $t_{M}$. To avoid large tax base losses from small MNEs, the non-haven therefore abandons its commitment and also undercuts the GMT for small MNEs. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In the Regime 4 equilibrium, both countries therefore split their tax rates and set their unconstrained tax rates for small MNEs, while adhering to the GMT for large MNEs.
Figure 2: Equilibrium tax rates for different GMT rates
@@
Notes: The figure shows the equilibrium tax rates in the non-haven (a) and the haven (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes (c). Parameter values given in Table 2 affect the precise levels of tax rates and regime-switching points, but do not affect qualitatively the sequence and pattern of tax-competition equilibria.

### v6

@@ 1.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its pre-tax profits $\pi$ from its non-haven headquarters to each tax haven affiliate $h$, and incurs transaction costs $C_{h}(\cdot)$ for doing so. These costs can be interpreted as concealment costs or as expected fines, if detected. We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. [^10] This specification is standard in the literature (e.g., Hines and Rice, , 1994; Huizinga and Laeven, , 2008; Suárez Serrato, , 2018; Bilicka et al., , 2024). The parameter $\delta$ captures the difficulty of profit shifting. [^11] This can be higher, for example, due to taxpayer audit (Lediga et al., , 2023).
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^12] The GMT tax base equals that of national corporate taxes, ignoring additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler, (2024).
@@
The optimal choice is
@@
which is independent of $\pi$ and therefore holds for all MNEs.
The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity. Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, independent of $\pi$. Equal tax base elasticities remove countries’ incentives for tax discrimination across MNEs in the absence of the GMT. This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^13] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al., (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes. Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark.
@@ 1.3 Governments’ tax setting choices
Before examining the effects of the GMT, we first solve the unconstrained optimisation problem in the absence of the GMT. The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, , 2023). For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues (e.g., Janeba and Schjelderup, , 2023; Hindriks and Nishimura, , 2025).
The government of the non-haven $n$ therefore maximises
$G_{n}$ $=\underbrace{\int_{\underline{\pi}}^{\infty}\widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)dF}_{\text{Private benefit}}+\lambda\underbrace{\int_{\underline{\pi}}^{\infty}t_{n}\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi dF}_{\text{Tax revenues}}$
@@
$\hskip 18.49988pt\hskip 18.49988pt+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ is the total true profits of MNEs.
Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
$G_{h}=\lambda\int_{\underline{\pi}}^{\infty}t_{h}\theta_{h}\pi dF=\lambda t_{h}\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi.$ (4)
@@
Solving the first-order conditions of the two sets of countries yields best responses: [^14] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
@@
Solving for the optimal tax rates in the unconstrained benchmark, which we call ‘Regime 0,’ gives
@@
We hereafter assume $\delta\in(0,\,3H/2)$ to ensure $t_{n}^{0}\in(0,1)$.
Two comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the MVPF, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the value of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. [^15] Note the difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
The share of shifted profits in Regime 0 is given by
@@
Hence, when the non-haven maximises tax revenues only ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium. For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
2 Theoretical results
@@ 2.1 Tax regimes
With a GMT in place, in the first stage of the game each country decides whether to commit to a single tax rate—equal to or above the GMT rate. Non-committing countries split rates in the second stage, setting different rates for large and small MNEs. Large MNEs with profits $\pi\in[\pi_{M},\infty)$ are subject to the GMT, while small MNEs with $\pi\in[\underline{\pi},\pi_{M})$ are not. Let $\phi$ denote the exogenous share of profits earned by large MNEs, i.e., the coverage rate of the GMT:
$\phi\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\int_{\underline{\pi}}^{\infty}\pi dF}\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\Pi}.$ (8)
Table 1. Possible Regimes of the Non-Cooperative Tax Game.
@@
Table 1 shows possible combinations of tax policies of the non-haven and the havens. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these regimes, if the non-haven adopts a single GMT rate, the havens set a uniform tax rate at the same level or higher and therefore earn no positive tax revenues. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven for small MNEs. Alternatively, if the havens set a single non-GMT or a single GMT rate, the non-haven never wants to split its tax rate, due to the constant tax base elasticity. Hence only five possible tax regimes remain.
Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier. As $t_{M}$ rises, the equilibrium moves through four additional regimes, characterised as follows.
@@
Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
@@
    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
@@
Proof: See Appendix 1.
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates.
@@
In Regime 1, when the GMT rate $t_{M}$ exceeds the havens’ unconstrained rate $t_{h}^{0}$, they must apply $t_{M}$ to large MNEs. The non-haven’s unconstrained rate remains above $t_{M}$, so it continues to set a single rate above $t_{M}$ for all MNEs. The havens then choose whether to apply $t_{M}$ also to small MNEs or use a split schedule. By committing to a single GMT rate, they can induce the non-haven to raise its uniform rate. This argument is strongest when the havens decide collectively, as assumed in our main model. [^16] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Supplemental Appendix B. In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens. The havens trade off gains from inducing a higher non-haven rate against revenue losses from not undercutting the non-haven’s rate even more for small MNEs. In Regime 1, the first effect dominates, so they commit to the GMT rate for all MNEs, leading all countries to set higher uniform tax rates than in Regime 0.
In Regime 2, with $t_{M}$ above the regime-switching rate $t_{M}^{1}$, the non-haven remains unconstrained and sets a single rate above $t_{M}$ for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. The higher $t_{M}$ reduces gains from a uniform rate, as the tax differential and thus shifted profits shrink. [^17] This is seen from Proposition 1(ii), where the tax differential $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases with $t_{M}$. At the same time, it increases losses from not competing with the non-haven for small MNEs. Consequently, the havens split their rates, applying $t_{M}$ to large MNEs and a lower rate to small MNEs.
@@
In Regime 3, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the non-haven can no longer set a rate above $t_{M}$ but is bound by it for large MNEs. By committing to $t_{M}$ for all MNEs, it induces the havens to raise their rates for small MNEs, analogous to the argument in Regime 1. In equilibrium, the non-haven applies $t_{M}$ to all firms, while the havens set $t_{M}/2$ for small MNEs.
@@
Finally, in Regime 4, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set $t_{M}$ for all MNEs. [^18] This is seen from Proposition 1(iii), where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases with $t_{M}$. To limit tax base losses from small MNEs, the non-haven abandons its commitment and lowers its tax rate for them. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In Regime 4, both countries split their rates, applying unconstrained rates to small MNEs and the GMT rate to large MNEs.
Fig.1. Equilibrium Tax Rates for Different GMT Rates.
@@
Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c). While the levels of tax rates and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.